OpenStudy (anonymous):

Phosphorus-32 has a half-life of 14.3 days. How many milligrams of phosphorus-32 remain after 71.5 days is you start with 4.00 mg of the isotope?

OpenStudy (anonymous):

Do you know the half life formula?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

N(t) = No(1/2)^(t/ t1/2) N(t) is the amount you have left after a certain time No is the initial mass of whatever that you have t is the time that passes t 1/2 is the half life time Think you can sub in the numbers from the question into this equation and evaluate?

OpenStudy (anonymous):

wow, no sorry, I never even learned this, don´t know why i´m being asked this.

OpenStudy (anonymous):

It says in the problem in pretty plain language what each are, like it says the half life of the phosphorous is 14.3 days, so thats t 1/2

OpenStudy (xishem):

In other words...\[N(t)=N_0 \frac{1}{2}^\frac{t}{t_{1/2}}\]\[N(71.5\ days)=4.00mg*0.5^\frac{71.5}{14.3}\]\[N(71.5\ days) = 0.125mg\]